# Chia-Hsiang Lin, , and Jhao-Ting Lin

### v1

@@ I Introduction
Most existing optical satellites acquire multispectral images (MSI) over diverse landscapes, while they often suffer from the so-called mixed pixel phenomenon as typical multispectral images in remote sensing have limited spatial resolution. This means that the spatial region of one pixel could cover a large area hence containing multiple substances, and thus a pixel could be considered as a mixture of several pure multispectral spectra of the underlying substances. To facilitate practical applications, we need to develop blind source separation (BSS) technology [1] to unambiguously recover the pure spectra from the mixed multispectral pixels for better classification and identification. To echo relevant BSS development in the hyperspectral remote sensing area [2], we term this BSS technology as multispectral unmixing (MU).
Different from hyperspectral unmixing (HU), which has more observations than sources, the MU problem mathematically corresponds to the underdetermined equation system, and is theoretically much more challenging than HU. In fact, the HU problem often induces NP-hard criteria, such as Craig’s geometry criterion [3, 4, 5] or the algebraic criterion called non-negative matrix factorization (NMF) [6]. Since MU corresponds to the even challenging underdetermined case, this is probably why it prevents researchers from tackling it, making the MU technique quite rarely seen in the literature. After conducting a detailed survey, we found one probably the most relevant work [7], where some seminal HU algorithms are extended for MU. Specifically, MU is attempted by the three HU algorithms, including Vertex Component Analysis (VCA) [8], N-FINDR [9], and NMF [10]. The three algorithms are selected partly because they (or their variations) are the most widely used unmixing methods, which are proven to be robust and reliable, and partly because they represent different HU categories showing unique advantages depending on the real scenarios [7]. It is interesting to remark that convex geometry plays a critical role in N-FINDR [9], which is based on the Winter’s criterion that attempts to find the largest simplex inscribed in the data convex hull [11]. This is converse to the aforementioned Craig’s criterion, which recovers the pure spectra using the vertices of the minimum simplex that contains all the data pixel points. Related HU theory, known as John ellipsoid criterion [12], has been proposed for high-performance HU to mitigate the NP-hardness of Craig’s criterion. John ellipsoid criterion can be implemented using convex optimization, with convex log-determinant objective function over the positive definite cone, and with the convex second-order cone constraints. This convex problem can be solved using the first-order optimizer within polynomial time, based on the fast iterative shrinkage-thresholding algorithm (FISTA) optimization [13].
@@
However, the pioneering MU work [7] does not consider the underdetermined scenarios wherein there are more sources than observations; for example, one of the experiments therein involves $P=8$ multispectral bands but only with $N=4$ sources. To understand that our work aims at solving the MU from a fundamentally different but more practical perspective, let us mathematically describe our problem as to recover the multispectral endmember matrix ${\bm{B}}\in\mathbb{R}^{P\times N}$ and the abundance matrix ${\bm{S}}\in\mathbb{R}^{N\times L}$ from the observable multispectral image ${\bm{Z}}_{m}={\bm{B}}{\bm{S}}\in\mathbb{R}^{P\times L}$ in the case of more sources than bands (i.e., $P<N$), where $P$ is the number of multispectral bands, $N$ is the number of sources/substances presented in ${\bm{Z}}_{m}$, and $L$ is the number of pixels. Note that what we are interested in is the more practical underdetermined MU case of $P<N$, because typical multispectral satellites often have less than 5 spectral bands, while a typical multispectral image often covers more than 5 substances. However, previous MU works all ignore the underdetermined issue, and merely consider scenarios with more bands than sources (i.e., $P>N$).
@@
With a fundamentally different approach, this work attempts to resolve the underdetermined issue by incorporating a virtual prism to further conduct the light-splitting function on the observed MSI. Inspired by our recent investigation on the spectral super-resolution technique [19], such a virtual light-splitting task is considered technically feasible. As a side remark, the spectral super-resolution in [19] is achieved by blending the advantages of both convex optimization (CO) and deep learning (DE), leading to the so-called CODE learning theory [20]. Let us get back to the required virtual prism. Since the light-splitting function is challenging, this paper tries to achieve so through the help of the very advanced quantum feature extraction techniques [21]. Specifically, we implement this prism function as a quantum deep network (QUEEN), and, to ensure that our overall MU algorithm is a blind one, we have to train the QUEEN in a fully unsupervised manner. Note that blind mechanism is often preferred in remote sensing [22]. In very recent remote sensing literature [23], QUEEN has been applied to well restore satellite-acquired hyperspectral images (HSI). As this technique, referred to as hyperspectral QUEEN (HyperQUEEN), has very solid mathematical ground (e.g., provable quantum full expressibility), its quantum network architecture will be employed in our MU method [23].
To make the above seemingly crazy idea more feasible, we propose to somewhat simplify the complexity of the light-splitting function. For example, unlike the method in [19] that performs about 10 times spectral super-resolution, the virtual prism we need basically just has to perform 2 times spectral super-resolution, because usually we have $2P\geq N$ in practice (e.g., a 4-band MSI covering 6 material substances). Here, we purposely emphasize that the prism is virtual (allowing us to fix the spectral response as a simple deterministic matrix [24]), so the prism-generated virtual HSI has no need to correspond to some real hyperspectral sensor. Therefore, it should be sufficient as long as the virtual HSI satisfies some fundamental properties of light splitting (e.g., non-negativity and spectral continuity), greatly simplifying the design of the prism architecture and thus increasing the technical feasibility of the proposed idea. With the above virtual quantum prism, the prism-generated virtual HSI is expected to possess some natural properties of HSI data, such as the desired simplex geometry structure [25]. This not only allows us to utilize the virtual HSI to judiciously formulate the MU problem, but also allows us to solve the problem by adopting convex geometry theory to effectively perform HU on the virtual HSI, followed by spectrally downsampling the extracted pure endmember spectra back to the multispectral domain to complete the MU task. Experimental evidences demonstrate the feasibility and superiority of our proposed MU algorithm.
The remaining parts of this paper are organized as below. In Section II, we illustrate our innovative prism-based idea for solving the underdetermined MU problem, and show how it induces the cycling data-fitting term in our proposed MU criterion. We also implement the prism using quantum deep learning, and accordingly develop an algorithm to realize the proposed criterion, where we plug a convex geometry mechanism to achieve efficient computing. As this is the first MU work that fundamentally considers the underdetermined scenario, we design an experimental protocol to evaluate the MU performance in Section III, where we demonstrate the superiority of our algorithm with ablation study. Finally, the conclusions are drawn in Section IV.
Some standard notations used in this paper are collectively presented hereinafter. $\mathbb{Z}_{++}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$, denote, respectively, the sets of positive integers, real-valued $n$-dimensional vectors, and real-valued $m$-by-$n$ matrices. The set difference is denoted as $\setminus$”. $\|\cdot\|_{F}$ and $\|\cdot\|_{1}$ denote, respectively, the Frobenius norm and the $\ell_{1}$-norm. The symbols $\bm{1}_{m}$ and $\bm{I}_{m}$ represent, respectively, an all-one column vector of dimension $m$ and an identity matrix of dimension $m$-by-$m$. The Kronecker product is defined as $\otimes$. Given a matrix ${\bm{M}}$, the notation ${\bm{M}}\geq\bm{0}$ means that ${\bm{M}}$ is elementwise non-negative; $[{\bm{M}}]_{(i,:)}$, $[{\bm{M}}]_{(:,j)}$, and $[{\bm{M}}]_{(i,j)}$ denote the $i$th row of ${\bm{M}}$, $j$th column of ${\bm{M}}$, and the $(i,j)$th entry of ${\bm{M}}$, respectively. $\textrm{DIAG}({\bm{Y}}_{1},\dots,{\bm{Y}}_{N})$ denotes the block-diagonal matrix with ${\bm{Y}}_{n}$ being the $n$th diagonal block for $n=1,\dots,N$ [11]. The convex hull, affine hull [11], and volume of a given set $\mathcal{S}$ are, respectively, represented as $\textrm{conv}\mathcal{S}$, $\textrm{aff}\mathcal{S}$, and $\textrm{volume}(\mathcal{S})$.
II The Proposed MU Algorithm
II-A Criterion Design
@@
where, considering the high complexity of the function that the virtual prism should achieve, we introduce the rather advanced quantum feature extraction technique [23] to learn unitary features for realizing the function $f$ in (II-A), as will be defined and detailed later.
@@
Though (II-A) forming the cycling in Figure 1 well resolves the underdetermined issue, the criterion itself is still ill-posed, and thus requires additional regularization. Below, we discuss the regularization of $({\bm{A}},{\bm{S}},{\bm{Z}}_{h},f)$. First, the regularization of the hyperspectral signature matrix ${\bm{A}}=[{\bm{a}}_{1},\dots,{\bm{a}}_{N}]$ is tricky. Accordingly to the well known Craig’s criterion in the remote sensing area [3], the endmember simplex $\textrm{conv}\{{\bm{a}}_{1},\dots,{\bm{a}}_{N}\}$ will hold a high resemblance to the minimum-volume data-enclosing simplex. In fact, under mild practical conditions on the data purity, Craig’s criterion has been proven to be a perfect conjecture with solid theoretical ground about two decades after it was proposed [29, 30]. This motivates us to regularize ${\bm{A}}$ using the volume of the endmemeber simplex, i.e.,
@@
where $\lambda>0$ is the regularization parameter empirically set as $\lambda:=0.1$, and the three non-negativity constraints are oriented from the natures of the respective quantities. As (II-A) is solved, we obtain the optimal $({\bm{A}}^{\star},{\bm{S}}^{\star})$, from which we have the final MU solution; specifically, we have the multispectral endmember matrix ${\bm{B}}^{\star}:={\bm{D}}{\bm{A}}^{\star}$ and the abundance matrix ${\bm{S}}^{\star}$ (cf. Algorithm 1).
II-B Geometry-Based Algorithm Design
Next, we design an algorithm to implement the proposed MU criterion (II-A). Directly handling it is too difficult even if we are simply talking about the optimization ordering of the four block variables. Thus, we propose to merge $({\bm{A}},{\bm{S}})$ into a single block variable, thereby allowing us to elegantly adopt convex geometry to simultaneously solve $({\bm{A}},{\bm{S}})$, as will be detailed later. Now, we just have three block variables, easier for us to decide the optimization ordering. Considering the ease of the initialization, we propose to first optimize $f$, followed by optimizing ${\bm{Z}}_{h}$, and finally the block variable $({\bm{A}},{\bm{S}})$ is optimized, as specified in Algorithm 1. We term our algorithm as prism-inspired multispectral endmember extraction (PRIME).
Algorithm 1 PRIME Algorithm for Solving (II-A)
@@ II-B Geometry-Based Algorithm Design
To initiate this alternating optimization scheme, we need to initialize ${\bm{Z}}_{h}$ and $({\bm{A}},{\bm{S}})$. For ${\bm{Z}}_{h}$, we propose to initialize it as a slightly perturbed light-splitting version of the $P$-band ${\bm{Z}}_{m}$, as detailed below. As discussed previously, the virtual ${\bm{Z}}_{h}$ has no need to correspond to some real HSI sensor, so we can flexibly assume that each band in ${\bm{Z}}_{m}$ is split into $\gamma\in\mathbb{Z}_{++}$ bands (i.e., $M=\gamma P$). Therefore, the spectral response matrix ${\bm{D}}$ is given by ${\bm{D}}\triangleq{\bm{I}}_{P}\otimes\bm{1}^{T}_{\gamma}\in\mathbb{R}^{P\times M}$. Considering that the concept of the $M$-band virtual ${\bm{Z}}_{h}$ is proposed to mitigate the underdetermined issue (thus requiring $M=\gamma P\geq N$), and that we usually have $2P\geq N$ in practice as discussed before, we can simply consider $\gamma:=2$ for conciseness. According to the physical meaning of light splitting, the $i$th MSI band $[{\bm{Z}}_{m}]_{(i,:)}$ can be split into two HSI bands, i.e., $0.5([{\bm{Z}}_{m}]_{(i,:)}\pm{[\bm{\theta}]}_{(i,:)})$ for some $[\bm{\theta}]_{(i,:)}$, which ensures the light-splitting relation of ${\bm{D}}{\bm{Z}}_{h}={\bm{Z}}_{m}$. Considering the desired continuity of the HSI bands, we can set $[\bm{\theta}]_{(i,:)}$ as $\frac{1}{4}([{\bm{Z}}_{m}]_{(i+1,:)}-[{\bm{Z}}_{m}]_{(i,:)})$ for uniform spectral sampling; further considering the boundary issue, the last term (i.e., the $P$th term) is set as $[\bm{\theta}]_{(P,:)}$ as $\frac{1}{4}([{\bm{Z}}_{m}]_{(P,:)}-[{\bm{Z}}_{m}]_{(P-1,:)})$. To be precise, the above generated $M$-band information $\widetilde{{\bm{Z}}}_{h}$ can be explicitly written as
It can be proven true that the $M$-band $\widetilde{{\bm{Z}}}_{h}$ is still of rank $P$, so it is slightly perturbed with Gaussian noise ${\bm{N}}$ whose energy is about $p:=5\%$ of $\widetilde{{\bm{Z}}}_{h}$, leading to the initialization of
${\bm{Z}}_{h}^{0}=\widetilde{{\bm{Z}}}_{h}+c{\bm{N}},~{}\textrm{with}~{}c\triangleq\sqrt{\frac{p\|\widetilde{{\bm{Z}}}_{h}\|_{F}^{2}}{\|{\bm{N}}\|_{F}^{2}}},$ (8)
The remaining task is to derive solutions for the three block variables in the MU criterion (II-A). According to (II-A), (5), (6), and (II-A), they amount to the following three subproblems, i.e.,
@@
which will be alternatively updated in Algorithm 1. To solve (II-B), we will design $f$ as a QUEEN-based quantum prism network, and train the network to learn the light-splitting relation between ${\bm{Z}}_{h}^{t}$ and ${\bm{Z}}_{m}$ using Adam optimizer [33]. As there is a long story behind the network design of $f$, it will be collectively presented in Section II-C, where (II-B) will be considered as the loss function during the training phase.
To solve (II-B), the non-negative constraint would make the optimization slow. Considering that the overall criterion (II-A) is non-convex, we may not need to pursue a global optimal solution. Thus, for efficient optimization, we solve the objective function in (II-B), followed by projecting the solution to the non-negative orthant to catch up with the constraint of (II-B), leading to the update of
@@
where $\Pi(\cdot)$ is the projector onto the non-negative orthant, as presented in Algorithm 1.
To solve (11), note that even without the regularizers $V({\bm{A}})$ and $\|{\bm{S}}\|_{1}$, the first term of (11) is already the NP-hard NMF problem. That said, we need to find a strategy to efficiently address (11). Our idea is to employ the convex geometry algorithm called Hyperplane-based Craig Simplex Identification (HyperCSI) [4], which is not only extremely fast but also exactly matches the physical meaning of (11), as illustrated below. Let us explain how the three terms in (11) associate with the mechanisms in HyperCSI. For the first term, the problem “$\min_{{\bm{A}},{\bm{S}}\geq\mathbf{0}}~{}\!\frac{\|{\bm{Z}}_{h}^{t+1}-{\bm{A}}{\bm{S}}\|^{2}_{F}}{\lambda}~{}\equiv~{}\min_{{\bm{A}},{\bm{S}}\geq\mathbf{0}}~{}\!\|{\bm{Z}}_{h}^{t+1}-{\bm{A}}{\bm{S}}\|^{2}_{F}$” (cf. $\lambda>0$) is exactly the unregularized NMF criterion for performing HU on ${\bm{Z}}_{h}^{t+1}$ [10, 4], and HyperCSI is exactly proposed for high-performance HU. For the second term $V({\bm{A}})$, it encourages a solution ${\bm{A}}=[{\bm{a}}_{1},\dots,{\bm{a}}_{N}]$ to form the minimum-volume simplex $\textrm{conv}\{{\bm{a}}_{1},\dots,{\bm{a}}_{N}\}$ (cf. (3)), and this is exactly what Craig’s criterion proposes to do [3]. This minimum-volume mechanism is exactly covered by HyperCSI, because HyperCSI is developed based on Craig’s minimum-volume HU criterion [4, Equation (8)]. For the third term $\|{\bm{S}}\|_{1}$, this is to promote the abundance sparsity of ${\bm{S}}$, as discussed in (4). Notice that in HyperCSI, instead of estimating the vertices of the minimum simplex, it equivalently estimates the hyperplanes forming the boundary of the minimum simplex. Using such a hyperplane geometry is exactly to capture the abundance sparsity of ${\bm{S}}$, as detailed in [4, Section III.E], where the key theory is that $[{\bm{S}}]_{(i,n)}=0$ (sparsity) if, and only if, the $n$th pixel of ${\bm{Z}}_{h}^{t+1}$ belongs to the $i$th hyperplane $\mathcal{H}_{i}\triangleq\textrm{aff}(\{{\bm{a}}_{1},\dots,{\bm{a}}_{N}\}\setminus\{{\bm{a}}_{i}\})$ [4, Section III.E]. Thus, identifying the boundary hyperplanes also promotes the abundance sparsity. Overall, both (11) and HyperCSI are to conduct HU using the minimum-volume and sparsity regularization mechanisms. Therefore, we update $({\bm{A}}^{t+1},{\bm{S}}^{t+1})$ of (11) by processing ${\bm{Z}}_{h}^{t+1}$ using HyperCSI [4], whose compression and radius parameters are set as $\eta:=1$ and $r:=1\textrm{E}-8$, respectively. We have completed the design of the PRIME algorithm for MU, which is summarized in Algorithm 1.

### v2

@@ I Introduction
Most existing optical satellites acquire multispectral images (MSIs) over diverse landscapes, while they often suffer from the so-called mixed pixel phenomenon, as typical MSIs in remote sensing have limited spatial resolution. This means that the spatial region of one pixel could cover a large area, hence containing multiple substances, and thus a pixel could be considered as a mixture of several pure multispectral spectra of the underlying substances. To facilitate practical applications, we need to develop blind source separation (BSS) technology [1] to unambiguously recover the pure spectra from the mixed multispectral pixels for better classification and identification. To echo relevant BSS development in the hyperspectral remote sensing area [2], we term this BSS technology as multispectral unmixing (MU).
@@
However, the pioneering MU work [3] does not consider the underdetermined scenarios wherein there are more sources than observations; for example, one of the experiments therein involves $P=8$ multispectral bands but only with $N=4$ sources. To understand that our work aims at solving the MU from a fundamentally different but more practical perspective, let us mathematically describe our problem as to recover the multispectral endmember matrix ${\bm{B}}\in\mathbb{R}^{P\times N}$ and the abundance matrix ${\bm{S}}\in\mathbb{R}^{N\times L}$ from the observable MSI ${\bm{Z}}_{m}={\bm{B}}{\bm{S}}\in\mathbb{R}^{P\times L}$ in the case of more sources than bands (i.e., $P<N$), where $P$ is the number of multispectral bands, $N$ is the number of sources/substances presented in ${\bm{Z}}_{m}$, and $L$ is the number of pixels. Note that what we are interested in is the more practical underdetermined MU case of $P<N$, because typical multispectral satellites often have less than $5$ spectral bands, while a typical MSI often covers more than $5$ substances. However, previous MU works all ignore the underdetermined issue, and merely consider scenarios with more bands than sources (i.e., $P>N$).
With a fundamentally different approach, this work attempts to resolve the underdetermined issue by incorporating a virtual prism to further conduct the light-splitting function on the observed MSI. Inspired by our recent investigation on the spectral super-resolution technique [4], such a virtual light-splitting task is considered technically feasible. As a side remark, the spectral super-resolution in [4] is achieved by blending the advantages of both convex optimization (CO) and deep learning (DE), leading to the so-called CODE learning theory [5]. Let us get back to the required virtual prism. Since the light-splitting function is challenging, this paper tries to achieve so through the help of the very advanced quantum feature extraction techniques [6]. Specifically, we implement this prism function as a quantum deep network (QUEEN), and, to ensure that our overall MU algorithm is a blind one, we have to train the QUEEN in a fully unsupervised manner. Note that blind mechanism is often preferred in remote sensing [7]. In very recent remote sensing literature [8], QUEEN has been applied to well restore satellite-acquired hyperspectral images (HSIs). As this technique, referred to as hyperspectral QUEEN (HyperQUEEN), has a very solid mathematical ground (e.g., provable quantum full expressibility), its quantum network architecture will be employed in our MU method [8].
To facilitate the above idea, a basic light-splitting function for reducing the complexity could already be sufficient for the MU application. Specifically, unlike the light-splitting method proposed in [4] achieving approximately $10$ times spectral super-resolution, the virtual prism in our MU method is set by default to perform $2$ times spectral super-resolution because usually we have $2P\geq N$ in practice (e.g., a $4$-band MSI covering $6$ material substances). However, as will be experimentally demonstrated, our MU method can be easily upscaled to perform higher order spectral super-resolution, if necessary (cf. Section IV-D). Here, we purposely emphasize that the prism is virtual (allowing us to fix the spectral response as a simple deterministic matrix [9]), so the prism-generated virtual HSI has no need to correspond to some real hyperspectral sensor. Therefore, it should be sufficient as long as the virtual HSI satisfies some fundamental properties of light splitting (e.g., non-negativity and spectral continuity), greatly simplifying the design of the prism architecture and thus increasing the technical feasibility of the proposed idea. With the above virtual quantum prism, the prism-generated virtual HSI is expected to possess some natural properties of HSI data, such as the desired simplex geometry structure [10]. This not only allows us to utilize the virtual HSI to judiciously formulate the MU problem, but also allows us to solve the problem by adopting convex geometry theory to effectively perform HU on the virtual HSI, followed by spectrally downsampling the extracted pure endmember spectra back to the multispectral domain to complete the MU task. Experimental evidences demonstrate the feasibility and superiority of our proposed MU algorithm.
The remaining parts of this paper are organized as below. In Section II, we review existing HU methods, highlighting their theoretical foundations and key advancements. Additionally, given the limited MU literature, we discuss relevant underdetermined BSS techniques to echo MU. In Section III, we illustrate our innovative prism-based idea for solving the underdetermined MU problem, and show how it induces the cycling data-fitting term in our proposed MU criterion. We also implement the prism using quantum deep learning, and accordingly develop an algorithm to realize the proposed criterion, where we plug a convex geometry mechanism to achieve efficient computing. As this is the first MU work that fundamentally considers the underdetermined scenario, we design an experimental protocol to evaluate the MU performance in Section IV, where we demonstrate the superiority of our algorithm with ablation study. Finally, the conclusions are drawn in Section V.
Some standard notations used in this paper are collectively presented hereinafter. $\mathbb{Z}_{++}$ denotes the sets of positive integers. $\mathbb{R}^{n}$ and $\mathbb{R}^{m\times n}$, denote, respectively, real-valued $n$-dimensional vectors and real-valued $m$-by-$n$ matrices. The set difference is denoted as $\setminus$”. $\|\cdot\|_{F}$ and $\|\cdot\|_{1}$ denote, respectively, the Frobenius norm and the $\ell_{1}$-norm. The symbols $\bm{1}_{m}$ and $\bm{I}_{m}$ represent, respectively, an all-one column vector of dimension $m$ and an identity matrix of dimension $m$-by-$m$. The Kronecker product is defined as $\otimes$. Given a matrix ${\bm{M}}$, the notation ${\bm{M}}\geq\bm{0}$ means that ${\bm{M}}$ is elementwise non-negative; $[{\bm{M}}]_{(i,:)}$, $[{\bm{M}}]_{(:,j)}$, and $[{\bm{M}}]_{(i,j)}$ denote the $i$th row of ${\bm{M}}$, $j$th column of ${\bm{M}}$, and the $(i,j)$th entry of ${\bm{M}}$, respectively. $\textrm{DIAG}({\bm{Y}}_{1},\dots,{\bm{Y}}_{N})$ denotes the block-diagonal matrix with ${\bm{Y}}_{n}$ being the $n$th diagonal block for $n=1,\dots,N$ [11]. The convex hull, affine hull [11], and volume of a given set $\mathcal{S}$ are, respectively, represented as $\textrm{conv}\mathcal{S}$, $\textrm{aff}\mathcal{S}$, and $\textrm{volume}(\mathcal{S})$.
II Related Works
In this section, we begin by reviewing existing HU methods, quite related to MU. With the increase in open spectral libraries, sparse regression (SA), a semi-supervised approach, has recently been proposed to address the HU problem, effectively reducing reliance on pure pixels. Sparse unmixing by variable splitting and augmented Lagrangian (SUnSAL) [12] and its constrained variant, C-SUnSAL, are classical methods based on the alternating direction method of multipliers (ADMM). Building on SUnSAL, SUnSAL-TV [13] introduces total variation regularization to impose spatial consistency. SUnSPI [14] incorporates spectral a priori information to promote sparsity in inactive signatures. Spatial filtering-based sparse unmixing (SFSU) [15] adopts the edge-preserving filtering strategy to address complex noise interference. With the rapid advancements in artificial intelligence (AI), SUnCNN [16] emerged as the first DE technique for sparse unmixing, utilizing a convolution neural network (CNN) architecture combined with a softmax activation layer to enforce the abundance sum-to-one constraint.
While the above approaches effectively enhance abundance estimation, geometry-based methods that operate in a fully blind manner (i.e., performing endmember extraction without much prior knowledge) are often preferred in remote sensing due to their broader applicability. Under the so-called pure-pixel assumption, vertex component analysis (VCA) [17] iteratively identifies the farthest pixel (i.e., one of the pure pixels) from the data centroid, then projects the data onto a direction orthogonal to the subspace spanned by the previously selected pixels. N-FINDR [18] employs Winter’s criterion to identify the largest simplex formed by the pure pixels within the data convex hull [11]. However, a significant challenge of the pure-pixel assumption lies in the mixed pixel phenomenon, which makes pure pixels sparse or absent, thereby limiting the effectiveness of these algorithms.
In contrast to Winter’s criterion, Craig’s criterion [19, 20, 21] recovers the pure spectra using the vertices of the minimum simplex that contains all the data pixel points. Non-negative matrix factorization (NMF) [22] is well-suited for HU due to its strong correlation with the low-rank nature of HSIs. By incorporating Craig’s criterion as a regularization term, minimum volume constrained NMF (MVC-NMF) [23] suppresses noise sensitivity and improves the accuracy of endmember and abundance estimation. Since Craig’s criterion and NMF often involve NP-hard criteria, John ellipsoid criterion [24] has been proposed as an alternative approach for high-performance HU, aiming to mitigate the NP-hardness. John ellipsoid criterion can be implemented using convex optimization, with convex log-determinant objective function over the positive definite cone, and with the convex second-order cone constraints. This convex problem can be solved using the first-order optimizer within polynomial time, based on the fast iterative shrinkage-thresholding algorithm (FISTA) [25].
Unlike HU, which typically involves more observations than sources, the MU problem mathematically corresponds to the underdetermined equation system and is theoretically much more challenging than HU. This heightened complexity likely prevents researchers from tackling MU, making it quite rarely seen in the literature. After conducting a detailed survey, we found one probably the most relevant work [3], where some seminal HU algorithms are extended for MU. Specifically, MU is attempted by the three HU algorithms, including VCA [17], N-FINDR [18], and NMF [26]. The three algorithms are selected partly because they (or their variations) are the most widely used unmixing methods, which are proven to be robust and reliable, and partly because they represent different HU categories showing unique advantages depending on the real scenarios [3].
@@ II Related Works
III The Proposed MU Algorithm
III-A Criterion Design
@@
where, considering the high complexity of the function that the virtual prism should achieve, we introduce the rather advanced quantum feature extraction technique [8] to learn unitary features for realizing the function $f$ in (III-A), as will be defined and detailed later.
@@
Though (III-A) forming the cycling in Figure 1 well resolves the underdetermined issue, the criterion itself is still ill-posed, and thus requires additional regularization. Below, we discuss the regularization of $({\bm{A}},{\bm{S}},{\bm{Z}}_{h},f)$. First, the regularization of the hyperspectral signature matrix ${\bm{A}}=[{\bm{a}}_{1},\dots,{\bm{a}}_{N}]$ is tricky. Accordingly to the well known Craig’s criterion in the remote sensing area [19], the endmember simplex $\textrm{conv}\{{\bm{a}}_{1},\dots,{\bm{a}}_{N}\}$ will hold a high resemblance to the minimum-volume data-enclosing simplex. In fact, under mild practical conditions on the data purity, Craig’s criterion has been proven to be a perfect conjecture with solid theoretical ground about two decades after it was proposed [35, 36]. This motivates us to regularize ${\bm{A}}$ using the volume of the endmemeber simplex, i.e.,
@@
where $\lambda>0$ is the regularization parameter empirically set as $\lambda:=0.1$, and the three non-negativity constraints are oriented from the natures of the respective quantities. As (III-A) is solved, we obtain the optimal $({\bm{A}}^{\star},{\bm{S}}^{\star})$, from which we have the final MU solution; specifically, we have the multispectral endmember matrix ${\bm{B}}^{\star}:={\bm{D}}{\bm{A}}^{\star}$ and the abundance matrix ${\bm{S}}^{\star}$ (cf. Algorithm 1).
III-B Geometry-Based Algorithm Design
Next, we design an algorithm to implement the proposed MU criterion (III-A). Directly handling it poses significant challenges, even if we are simply considering the ordering of the four block variables under the alternating optimization scheme. Thus, we propose to merge $({\bm{A}},{\bm{S}})$ into a single block variable, thereby allowing us to elegantly adopt convex geometry to simultaneously solve $({\bm{A}},{\bm{S}})$, as will be detailed later. Now, we just have three block variables, easier for us to decide the optimization ordering. Considering the ease of the initialization, we propose to first optimize $f$, followed by optimizing ${\bm{Z}}_{h}$, and finally the block variable $({\bm{A}},{\bm{S}})$ is optimized, as specified in Algorithm 1. We term our algorithm as prism-inspired multispectral endmember extraction (PRIME).
Algorithm 1 PRIME Algorithm for Solving (III-A)
@@ III-B Geometry-Based Algorithm Design
With the light-splitting-driven algorithm initialization detailed in Appendix -A, we now focus on deriving the solutions for the three block variables in the MU criterion (III-A). According to (III-A), (5), (6), and (III-A), they amount to the following three subproblems, i.e.,
@@
which will be alternatively updated in Algorithm 1. To solve (III-B), we design $f$ as a QUEEN-based quantum prism network, and train the network to learn the light-splitting relation between ${\bm{Z}}_{h}^{t}$ and ${\bm{Z}}_{m}$ using adaptive moment estimation (Adam) optimizer [38]. The detailed rationale behind the design of $f$ will be thoroughly discussed in Section III-C, where the objective function in (III-B) will elegantly serve as the loss function during the model training procedure.
To solve (III-B), the non-negative constraint would make the optimization slow. Considering that the overall criterion (III-A) is non-convex, we may not need to pursue a global optimal solution. Thus, for efficient optimization, we solve the objective function in (III-B), followed by projecting the solution to the non-negative orthant to catch up with the constraint of (III-B), leading to the update of
@@
where $\Pi(\cdot)$ is the projector onto the non-negative orthant.
To solve (10), note that even without the regularizers $V({\bm{A}})$ and $\|{\bm{S}}\|_{1}$, the first term of (10) is already the NP-hard NMF problem. That said, we need to find a strategy to efficiently address (10). Our idea is to employ the convex geometry algorithm called Hyperplane-based Craig Simplex Identification (HyperCSI) [20], which is not only extremely fast but also exactly matches the physical meaning of (10), as illustrated below. Let us explain how the three terms in (10) associate with the mechanisms in HyperCSI. For the first term, the problem “$\min_{{\bm{A}},{\bm{S}}\geq\mathbf{0}}~{}\!\frac{\|{\bm{Z}}_{h}^{t+1}-{\bm{A}}{\bm{S}}\|^{2}_{F}}{\lambda}~{}\equiv~{}\min_{{\bm{A}},{\bm{S}}\geq\mathbf{0}}~{}\!\|{\bm{Z}}_{h}^{t+1}-{\bm{A}}{\bm{S}}\|^{2}_{F}$” (cf. $\lambda>0$) is exactly the unregularized NMF criterion for performing HU on ${\bm{Z}}_{h}^{t+1}$ [26, 20], and HyperCSI is exactly proposed for high-performance HU. For the second term $V({\bm{A}})$, it encourages a solution ${\bm{A}}=[{\bm{a}}_{1},\dots,{\bm{a}}_{N}]$ to form the minimum-volume simplex $\textrm{conv}\{{\bm{a}}_{1},\dots,{\bm{a}}_{N}\}$ (cf. (3)), and this is exactly what Craig’s criterion proposes to do [19]. This minimum-volume mechanism is exactly covered by HyperCSI, because HyperCSI is developed based on Craig’s minimum-volume HU criterion [20, Equation (8)]. For the third term $\|{\bm{S}}\|_{1}$, this is to promote the abundance sparsity of ${\bm{S}}$, as discussed in (4). Notice that in HyperCSI, instead of estimating the vertices of the minimum simplex, it equivalently estimates the hyperplanes forming the boundary of the minimum simplex. Using such a hyperplane geometry is exactly to capture the abundance sparsity of ${\bm{S}}$, as detailed in [20, Section III-E], where the key theory is that $[{\bm{S}}]_{(i,n)}=0$ (sparsity) if, and only if, the $n$th pixel of ${\bm{Z}}_{h}^{t+1}$ belongs to the $i$th hyperplane $\mathcal{H}_{i}\triangleq\textrm{aff}(\{{\bm{a}}_{1},\dots,{\bm{a}}_{N}\}\setminus\{{\bm{a}}_{i}\})$ [20, Section III-E]. Thus, identifying the boundary hyperplanes also promotes the abundance sparsity. Overall, both (10) and HyperCSI are to conduct HU using the minimum-volume and sparsity regularization mechanisms. Therefore, we update $({\bm{A}}^{t+1},{\bm{S}}^{t+1})$ of (10) by processing ${\bm{Z}}_{h}^{t+1}$ using HyperCSI [20], whose compression and radius parameters are set as $\eta:=1$ and $r:=1\textrm{E}-8$, respectively. We have completed the design of the PRIME algorithm for MU, which is summarized in Algorithm 1.